GreenHab Report – 03 May 2018

Date: 3May2018

Environmental control: Ambient with door open in the afternoon

Shade cloth: on

Fan: Briefly

Average temperatures: 39.0℃

Low: 15.2℃

High: 43.5℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 8 Gallons

Time(s) of watering for crops: Twice: at 0915 and again at 1615

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: The plants are very happy with their care, but are eager to try some of this ‘Starbucks’ they’ve been hearing so much about, so if you would please pass some this way…

Support/supplies needed: None

EVA Report – 03 May 2018

Sol 11 Crew 193 EVA Summary
Earth-Date 3MAY2018

EVAs _
There were no EVAs planned today, as we had visitors this morning that lasted into the afternoon. After they left we had some rain, which precluded any contingency EVA that we had planned.

We did have an Engineering EVA this morning, in which our CO used an ATV to make contact with the visitors from another planet.

Journalist Report May 2nd

Journalist Report – Earth-Date 02May2018 / Sol 10

Author: Eric Shear

This has been a day of cleaning and humor. In light of both EVAs being cancelled today due to extreme muddy conditions, we have made the best of a bad situation and channeled our energies into preparations for the school field trip tomorrow. We have cleaned the hab to the best of our ability, kept up with our experiments, and made plans for receiving the kids.

Our astronomer took advantage of a brief peek of the sun through the clouds to photograph one small prominence on the solar limb. The general conclusion he reached from these observations is that the sun is undergoing a quiescent period of activity, which is extremely fortunate for our mission.

We learned some more ASL from our deaf crew member. Some of us managed to guess the signs correctly before he taught them to us!

We are looking forward to the end of our rotation, and have begun our final reports in anticipation of that big day.

Sol Summary Report – May 2nd

Crew 193 Sol Summary Report – 1MAY2018

Sol 9

Summary Title: Can’t go out cause Sol this mud.

Author: David Attig, Crew Engineer

Mission Status: Nominally muddy

Sol Activity Summary: The day started off while preparing to go out on EVA 18. Unfortunately, this was called off due to the mud outside from the rain (on Mars?) overnight. The afternoon EVA was also cancelled. The crew spent most of the day working on personal projects and relaxing. We practiced ASL some more, and enjoyed a Martian feast.

Look Ahead Plan: Tomorrow we have a group of children coming to visit us on Mars. We look forward to enlightening them on the happenings up here, and what we need to survive.

Anomalies in work:

1. Muddy mud, muddying up the muck.

Weather: Overcast and muddy in the AM, with off-and-on rain. The sun came out in the afternoon, until we were under threat of a thunderstorm (which appears to be passing us).

Crew Physical Status: Feelin Fine.

EVA: Morning EVA was cancelled today, due to inclement weather (too muddy). Afternoon EVA cancelled for same reason.

Reports filed:

GreenHab Report

Astronomy Report

EVA Request

Science Report

Sol Summary

Operations Report

EVA Report

HSO Report

Reports to be filed:

Journalist Report

Photo Report & Photo of the Day

Support Requested:

1. A Starbucks franchise on Mars.

Operations Report – May 2nd

Crew 193 Operations Report 02May2018

SOL: 10

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 7:00 pm to 7:55 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 220 gallons.

Water (static) – 290 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134505.5

Toilet tank emptied: yes
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Muddy

Summary of internet: We’re out, same as normal.

Summary of suits and radios: One piece suit #10 is repaired. Still work to do on two piece suits. One suit charger repaired, repair of other suit chargers still in progress.

Summary of Hab operations: Muddily nominal

Summary of GreenHab operations: Muddy

Summary of ScienceDome operations: Muddy

Summary of RAM operations: None today, Muddy

Summary of health and safety issues: Muddy

Questions, concerns and requests to Mission Support: Less mud. And an escalator.

Astronomy Report – May 2nd

Astronomy Report

Name: Eric Shear

Crew: 193
Date: May 2, 2018

Sky Conditions: Heavy clouds & rain for most of the day. The sun came out between the clouds in the afternoon.

Wind Conditions: Low

Observation Start Time: 1430

Observation End Time: 1500


On eye-viewing, the sun appeared to have no obvious features. The photo taken below has a tiny prominence.

Photos were taken with the Skyris camera and the best images processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached.

Objects Viewed: Sun

Problems Encountered:

There was a light drizzle despite the sun being out among the clouds, so I waited 10-15 minutes before it stopped before taking the blanket off the telescope.

Green Hab Report – May 2nd

Hello Mission Control,

Please find below the Green Hab Report for May 2, 2018.

Green Hab Report

Date: 2 May 2018

Environmental control:

Ambient with door open but only for few hours in the afternoon

Shade cloth: on

Fan: not used today

Average temperatures: 34.5 C

Low: 15.4℃

High: 52.1℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 10 Gallons

Time(s) of watering for crops: Twice: In the morning at 1000am and afternoon 445pm.

Morning research observations: None

Changes to research plants: None

Aquaponics: None

Narrative: No significant changes.

Support/supplies needed: None

Science Report May 2nd

Hello mission control,

Here is the Science Report for 2 May 2018

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs occurred today due to weather.

2. Sunspot and Solar Flare Monitoring: a very small prominence was found on the sun today.

3. Shortwave Texting on EVA’s: No EVAs occurred today due to weather

4. GPS Route Measurements: No EVAs occurred today due to weather

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: No EVAs occurred today due to weather




Crew 193 XO and Science Officer

EVA Report May 2nd

Crew 193 EVA Report 02May2018

Sol 10 Crew 193 EVA 18 & 19 Summary

Earth-Date 2MAY2018

EVAs 18 & 19

Both EVAs 18 and 19 were cancelled due to extremely wet and muddy conditions on Mars. No science or exploration objectives were achieved today. No ATVs or rovers were used. No Gold crew members ventured outside the Hab on EVA.

Copyright © The Mars Society. All rights reserved. | Main Site